The equilibrium constant, Kp, for the following reaction is 0.160 at 298 K. 

Calculate K, for this reaction at this temperature. 

2NOBr(g)  2NO(g) + Br2(g) 

Kc =
Answer #1

Given reaction & NOB019) a NO(g) + B x (9) Kp = 0.16 at da8K. we know that I K = Kc CRYJAR An = [number of moles of gaseous pKc = . Kp (RT)An Kc = 0.16 (8-314x298) * = 0.16 - 2477-572 = 6.4579X1os = HE = 6. 46x105